OpenStudy (anonymous):

Another Simplifying Equation: this time there is no perfect squares except ..... 12z^2+25z+12 / 3z^2-2z-8 I know how to simplify the denominator which is (3x+4) and (x-2) but I do not know how to simplify the numerator. (method: quadratic equation??)

OpenStudy (souvik):

12z^2+25z+12 =12z^2+16z+9z+12 =4z(3z+4)+3(3z+4) =(3z+4)(4z+3)

OpenStudy (anonymous):

YES thank you! I got it but! , (how did you found out from 25z = 16z+9z??) What is the first step to simplify this equation??

OpenStudy (souvik):

you have (3x+4) in denominator ..if i get (3x+4)in numerator ...i can cancel this out..

OpenStudy (anonymous):

first find two numbers whose sum is 25 and product 12.12=144 which are 16 and 9

OpenStudy (souvik):

right..@Curious9876

OpenStudy (anonymous):

Nw break the expression 25z as 16z+9z 12z^2+16z+9z+12=4z(3z+4)+3(3z+4)=(4z+3)(3z+4)

OpenStudy (anonymous):

Yup @souvik